QUESTION

(a). Suppose that Beet Root Company has a 10 year preferred stock issue that pays a 15% dividend. The par value of each share is K80. The stocks are currently trading for K85. The going rate of interest in the market is 12%. What is the price of these shares?


(b). ABC Company is experiencing a period of rapid growth. Earnings and dividends

are expected to grow at a rate of 18 percent during each of the next two years, 15

percent in the third year, and at a constant rate of 6 percent annually thereafter.

ABC’s last dividend, which has just been paid, was K1.15. If the required rate of

return on the stock is 12 percent, what is the price of the stock today?


(c). What is the expected price of XYZ’s stock today if XYZ company’s dividend in a year’s time is expected to be K6.00 per share and grow at 12 percent for another 4 years, then 6 percent thereafter? XYZ shareholders require a return of 10 percent on the stock.  


1
Expert's answer
2021-03-14T19:52:38-0400

(a) find the formula:

"P=\frac{D}{(1+i)^k}+\frac{Pm}{(1+i)^n}"

"D=0.15\times(80)= 12"

k=1

n=15

Pm=80

"P=\frac{D}{(1+i)^k}+\frac{Pm}{(1+i)^n}=\frac{12}{(1+0.12)^1}+\frac{85}{(1+0.12)^{15}}=26.24"


(b)

find the formula:

"P=\frac{D(1+g)^n}{(1+i)^n}+\frac{D(1+g)^n}{(1+i)^n}+{D(i-g)}\times\frac{1}{(1+i)^5}=\frac{1.15(1+0.18)^2}{(1+0.12)^2}+\frac{1.15(1+0.15)^3}{(1+0.12)^3}+{1.15(0.12-0.06)}\times\frac{1}{(1+0.12)^5}=1.28+1.24+0.04=2.56"


(c) find the formula:

"P=\frac{D(1+g)^n}{(1+i)^n}+D(i-g)\times\frac{1}{(1+i)^5}=\frac{6(1+0.12)^4}{(1+0.10)^4}+6(0.1-0.06)\times\frac{1}{(1+0.1)^5}=6.60"